(Please show work)

Mcintarfer Corporation issued $200,000 of 8​%, 15​-year bonds payable on January​ 1, 2019. The market interest rate when the bonds were issued was 11​%. Interest is paid semiannually on January 1 and July 1. The first interest payment is July​ 1, 2019. Using the​ effective-interest amortization method, how much interest expense will Mcintarfer record on July​ 1, 2019​? Use Excel®  to calculate the issue price.

            $8,602
Answer #1

Answer:

Answer is Interest expense for July 1 ,2019 is $8,602
Bonds Par Value = $200,000
Interest Rate (semi annual ) =8% * 1/2 =4%
Market Rate semi annual = 11% * 1/2 =5.50%
Number of Periods = 15 Years = 30 Periods
Interest Payable semi annual = $200,000 * 4% = $8,000
issued price of Bonds equal to all future payments discounted at market rate
Issued Price = Interest Payment * Present value Annuity Factor at 5.50% 30 periods + Par value * PV factor at 5.50% at 30 th period
Issue price of Bonds = $8,000 * 14.53375   + $200,000 * 0.20064
Issue price of Bonds = $116,270 +$ 40,128 = $156,398
Interest expense = Issued price * Market rate of Interest = $156,398 * 5.50% =$8,601.89 =$8,602
Please rate my answer...in case any doubt, post a comment and I will try to resolve the doubt ASAP…thank you